If E(X) 5, V(X) = 3, and Y XX, what is the variance of Y? 5
Answer #1

Solution :

Var(Y) = 1/9 V(X) + V(X)

= 1/9 * 3 + 3

= 1/3 + 3

= 10 / 3

Variance of Y = 10/3 = 3.333